OpenStudy (anonymous):

Solve for x. 12x - 8 = 8x - 20 x = -3 x = -1 x = 1 3 x = 3

OpenStudy (astrophysics):

Start off by collecting like terms

OpenStudy (anonymous):

12x - 8 = 8x - 20 -8x -8x

OpenStudy (astrophysics):

\[12x-8x=-20+8\] now solve for x

OpenStudy (misty1212):

lol start by trying one yourself, this is like the fifth identical question

OpenStudy (anonymous):

lol its about time you guys let me try i fel likea baby i thx hol up

OpenStudy (anonymous):

okay, XD

OpenStudy (astrophysics):

You should try them out before posting it

OpenStudy (astrophysics):

And then if you're stuck just ask for help

OpenStudy (misty1212):

also @Astrophysics there are no like terms to collect in \[12x-8=8x-20\]

OpenStudy (anonymous):

4-8=20

OpenStudy (anonymous):

4x-8=-20

OpenStudy (astrophysics):

Like terms...x's on one side and the integers on the other.

OpenStudy (anonymous):

don't forgot the x

OpenStudy (anonymous):

and the negative sign

OpenStudy (anonymous):

4x-8=-20 +8 +8

OpenStudy (anonymous):

4x=-12

OpenStudy (anonymous):

I think you can do the rest....

OpenStudy (anonymous):

divide both sides by -4

OpenStudy (anonymous):

3

OpenStudy (astrophysics):

-3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

your dividing both sides by -4, not 4

OpenStudy (astrophysics):

No, you're dividing both side by 4

OpenStudy (astrophysics):

the 12 is just a negative